I have two problems I need help setting up. hers the first one.
Jerry is a purchasing agent for his company,and has a budget of $750,000 to purchase 600 new computers for the company staff. He has narrowed his choices to two different brands: Lateway, with a cost of 1,200 per computer, and cracintosh, with a cost of $1,450 each. In order to spend the budgeted amount exactly, how many employees will each receive a Cranintosh computer?
second problem
Jerry and his girlfriend Gerri live 326 miles apart. One day, Jerry decides to surprise Geri by driving to her place unannounced. He leaves at 10:00 a.m. and drives at an average speed of 68 m.p.h. Of course, it just so happens that Gerri had the same idea, and she left her place to surprise Jerry at 9:30 a.m., driving at an average speed of 76 m.p.h. At what time do they pass each other on the road? How far are they from Jerrys' place at that momement?
third problem
Jerry will soon attend a sold-out concert at a local amphitheater. There are two types of seating available: reserved seating close to the stage and general admission seating on the lawn. Reserved tickets are $34 and lawn tickets are $20. the total seating in the amphitheater is 16,000. If the total gate receipts (profit from ticket sales) are $409,600 how many folks will join Jerry on the lawn?

Answer by stanbon(48568)

Jerry is a purchasing agent for his company,and has a budget of $750,000 to purchase 600 new computers for the company staff. He has narrowed his choices to two different brands: Lateway, with a cost of 1,200 per computer, and cracintosh, with a cost of $1,450 each. In order to spend the budgeted amount exactly, how many employees will each receive a Cranintosh computer?
---------------------------------
Value + value = value
1200x + 1450(600-x) = 750,000
1200x + 1450*600 - 1450x = 750,000
-250x = -120000
x = 480 (# of employees will receive a Lateway)
------
600-x = 120 (# of employees will receive a Cranintosh)
---------------

second problem
Jerry and his girlfriend Gerri live 326 miles apart. One day, Jerry decides to surprise Geri by driving to her place unannounced. He leaves at 10:00 a.m. and drives at an average speed of 68 m.p.h. Of course, it just so happens that Gerri had the same idea, and she left her place to surprise Jerry at 9:30 a.m., driving at an average speed of 76 m.p.h. At what time do they pass each other on the road? How far are they from Jerrys' place at that momement?
----
J to G DATA:
rate = 68 mph ; time = x hrs ; distance = 68x miles
----------------
G to J DATA:
rate = 76 mph ; time = (x+1/2) hrs ; distance = 76(x+(1/2)) miles
----------------------
Equation:
distance + distance = 326 miles
68x + 76(x+(1/2)) = 326
68x + 76x + 38 = 326
144x = 288
x = 2 hrs (They pass each other after Jerry has driven 2 hrs; at 12:00
---
At that time Jerry is 2*68 = 136 miles from his starting point.
=====================================================================
third problem
Jerry will soon attend a sold-out concert at a local amphitheater. There are two types of seating available: reserved seating close to the stage and general admission seating on the lawn. Reserved tickets are $34 and lawn tickets are $20. the total seating in the amphitheater is 16,000. If the total gate receipts (profit from ticket sales) are $409,600 how many folks will join Jerry on the lawn?
---------------------
Equation:
Quantity Eq::R + L = 16000
Value Eq::::34R + 20L = 409,600
--------------------
Multiply through the Quantity Eq. by 34 to get:
34R + 34L = 34*16000
---
Subtract that from the Value Eq to get:
-14L = -134400
Divide to solve for L:
L = 9600 (# that will join Jerry on the lawn)
=================================================
Cheers,
Stan H.